y"=2y' reducing this into y


1
Expert's answer
2021-06-29T04:39:54-0400

Given the equation y"=2y'

Rewrite the equation to have

y"-2y'+0y=0

Where

p=-2

q=0

First of all,we should find the roots of the characteristics equation

"q+(k^2+kp)=0"

"0+(k^2-2k)=0"

"\implies k(k-2)=0"

"\therefore k_1=0"

"k_2=2"

The general solution is of the form

"y=C_1e^{k_1x}+C_2e^{k_2x}"

"y=C_1e^{0x}+C_2e^{2x}"

"\therefore y=C_1+C_2e^{2x}"